OpenStudy (anonymous):

Will fan and medal. Simplify the expression: 6x ---- 9(x+y) I know it's 2 --- 3 but I don't know how to do the variables exactly

OpenStudy (anonymous):

@Photon336

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (anonymous):

\[\frac{ 6x }{9(x+y) }\]

OpenStudy (photon336):

I get it now you need to simplify that expression

OpenStudy (photon336):

well, to simpify this you would need to divide both numbers by the greatest common factor. we see that we have 6 and 9 here so our greatest common factor would be 3 \[\frac{ 6x }{ 9(x+y) } = \frac{ 6/3(x) }{ 9/3(x+y) } = \frac{ 2x }{ 3(x+y) }\]

OpenStudy (anonymous):

Yes. Wait so the variables don't change in equations like this?

OpenStudy (photon336):

no because we cant factor out an x can you see why?

OpenStudy (anonymous):

I think so... because it's in the parenthesis right?

OpenStudy (photon336):

Think of it this way, whenever we have two variables added together like this (x+y) we cant factor out an x because an x is not in every term. say if we had \[x^{2}+xy = x(x+y)\] we could factor out an x because an x is in every term. but in our problem we can't factor out anything. \[(x+y) \]